OpenStudy (anonymous):

Find the 6th term of the sequence with t1 = -4 and tn = 5^tn-1

OpenStudy (anonymous):

t1 = -4 and tn = 5t^n-1 *

Parth (parthkohli):

You just substitute.

OpenStudy (anonymous):

-4 -1?

Parth (parthkohli):

Can you latexify it please?

OpenStudy (anonymous):

huh?

Parth (parthkohli):

Can you write the question with equation editor?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[t ^{1} and t ^{n} = ^5{t} n-1\]

OpenStudy (anonymous):

5t is squared

OpenStudy (anonymous):

not only 5

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow t_n = 5^{t - 1} }\) This?

OpenStudy (anonymous):

yeah but with n?

Parth (parthkohli):

Ok

Parth (parthkohli):

I get it :D

OpenStudy (anonymous):

ok :)

Parth (parthkohli):

So if n is 2 and (n - 1) is 1, then we can say: \(\Large \color{Black}{\Rightarrow t_2 = 5^{(t_1)} }\) \(\Large \color{Black}{\Rightarrow t_2= 5^{-4} }\) Just keep doing like this.

OpenStudy (anonymous):

ok

Parth (parthkohli):

I think you meant 5 * (tn -1)

Parth (parthkohli):

Am I right?

OpenStudy (anonymous):

yes

Parth (parthkohli):

So then it's a geometric sequence. \(\Large \color{Black}{\Rightarrow -4 \times 5^{n - 1} }\) Here \(\Large \color{Black}{\Rightarrow -4 \times 5^{6 - 1} }\) \(\Large \color{Black}{\Rightarrow -4 \times 5^5 }\) \(\Large \color{Black}{\Rightarrow -4 * 3125 }\) \(\Large \color{Black}{\Rightarrow -12500 }\)